OpenStudy (anonymous):

If f(x)=(1/3)x^3, then find k so that f(k)-f(3)=0

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

So you plug in K to the equation and 3 to the equation, and subtract?

OpenStudy (amistre64):

when does subtraction equate to zero?

OpenStudy (anonymous):

When both subtracting terms are the same.

OpenStudy (anonymous):

or equivalent.

OpenStudy (amistre64):

a-b=0, if a=b yes

OpenStudy (amistre64):

so f(k) has to be equal to f(3), what is k?

OpenStudy (anonymous):

So i just plug in a 3 and the answer will be the value of k?

OpenStudy (amistre64):

f(k) - f(3) = 0, implies that f(k) = f(3) yes

OpenStudy (amistre64):

1/3 (k^3) - 1/3 (3^3) = 0 1/3 (k^3 - 3^3) = 0 k^3 - 3^3 = 0 k^3 = 3^3

OpenStudy (anonymous):

so K=3

OpenStudy (amistre64):

yep ...

OpenStudy (anonymous):

Haha, okay Thanks!

OpenStudy (amistre64):

yw